economic order quantity

  1. Assuming that there is no safety stock and that the present stock level is 400 components, when should the next order be placed? (Assume a 360-day year.) (1 mark)

    COVID-19 Ltd produces a specialized product; LOCKDOWN which has a constant monthly demand of 4000 units. The LOCKDOWN requires a component which COVID-19 Ltd purchases from a supplier at GHS10 per unit. The component requires a three-day lead time from the date of order to the date of delivery. The ordering cost is GHS0.60 per order and the holiday cost is 10% per annum.

    Required:

  2. Calculate

  3. The economic order quantity.                                                              (1.5 marks)

  4. The number of orders required per year.                                              (1 mark)

  5. The total cost of ordering and holding the components for the year.    (1 mark)

  6. Assuming that there is no safety stock and that the present stock level is 400 components, when should the next order be placed? (Assume a 360-day year.) (1 mark)
Answer #1

Holding Cost Economic ordering quantity (E08) 2x Annual Demand Ordering Cost/order r/unit Annual Demand = 4000 x 12 = 48000 uAverage Daily Usage Total Holo iding Cost - Annual Demand X Holding Cost/ unit 48000 x 1 GHS 48000 Reorder point = Lead time
Answer #2

Given, Monthly Demand = 4000 units, Annual Production = 48000 units, Purchase price = GHS \(10 /\) unit, Ordering cost= =GHS 0.60 per order, Carrying cost \(=10 \%\) per annum \(=\mathrm{GHS} 10^{*} 10 \%=\mathrm{GHS} 1\)

ECONOMIC ORDER QUANTITY

$$ \begin{array}{r} \mathrm{EOQ}=\mathrm{V} 2 \mathrm{AO} / \mathrm{C} \\ \begin{array}{r} =\sqrt{2*48000*0.60} / 1 \\ =\sqrt{57600}  \\ =240 UNITS \end{array} \end{array} $$


NUMBER OF ORDERS PER YEAR

Total Annual Orders/EOQ

=48000 / 240 

=200 Orders


Total Ordering cost:

Annual ordering cost = no. of orders placed in a year \(x\) cost per order

$$ 200 \text { orders }^{*} \mathrm{GHS} 0.60=\mathrm{GHS} 120 $$

Total Handling cost:

Annual holding cost = Order quantity/2 x holding cost per unit per year

$$ 240 / 2 \text { units }^{*} \mathrm{GHS} 10^{*} 10 \%=\mathrm{GHS} 120 $$

NEXT ORDER BE PLACED Reorder point = daily demand x lead time + safety stock

Lead Time is not given, \(=(48000 / 360)=133\) units 

Already have 400 units.

Per Order takes 1.8 days, there is no need for the order to be placed now. Order placed after 240 units.


answered by: studyres